# Cavity-induced quantum droplets

### v1

@@ Abstract
The four modes of $\mathcal{K}_{\rm C}$ form a set of mutually interacting quantum harmonic oscillators with eigenfrequencies $\{\omega_{\bm{k}},\omega_{\bm{k}},\omega_{\bm{k}},\Omega\}$ [39], where
$\Omega=\sqrt{\omega_{\bm{k}}^{2}+16\frac{k^{2}}{M}V_{\rm C}(\bm{k},\bm{k})n}\,$ (5)
denotes the frequency of the roton mode. It involves via Eq. (3) the Fourier transform of the cavity-mediated interaction potential
$V_{\rm C}(\bm{p},\bm{p}^{\prime})=\frac{\mathcal{I}V}{32}\left[\frac{\sqrt{\pi}\xi}{L}\operatorname{erf}\left(\frac{L}{2\xi}\right)\right]^{4}\sum_{\bm{k},\bm{k}^{\prime}\in\mathcal{K}_{C}}\delta_{\bm{p}\bm{k}}\delta_{\bm{p}^{\prime}\bm{k}^{\prime}}\,.$ (6)
The roton softness is determined by the effective cavity-induced interaction strength $\mathcal{I}$. It can be tuned by changing the pump strength $h_{0}$ or the cavity detuning $\Delta_{\rm C}$, but can also be fine-tuned by adopting any other cavity feature. The roton mode is responsible for the leading-order energy contribution
$E_{\text{ac}}=\frac{1}{2}\left(\Omega-\omega_{\bm{k}}\right)\,$ (7)
of the cavity quantum fluctuations [39], which corresponds to the difference between the respective energies with and without coupling to the cavity and vanishes in the limit $\mathcal{I}\to 0$. Combining Eq. (4) with Eq. (7) yields the effective beyond-mean-field energy $E_{0}=E_{\rm mf}+E_{\text{ac}}$ of the condensate ground state. The chemical potential then follows as $\mu_{0}=gn+(\partial E_{\text{ac}}/\partial N)_{V}$.
Results.
Figure 1: Effective ground-state energy per atom $E_{0}/N$ with its mean-field ($E_{\text{mf}}$) and cavity-induced ($E_{\text{ac}}$) quantum fluctuation contributions plotted against the inverse cloud size $L$ for fixed atom number $N=10^{3}$. The vertical gray dashed line marks the energy minimum corresponding to the equilibrium droplet size $L_{0}$. The red dashed line indicates $E_{\text{ac}}^{(\infty)}$ for an infinite-range cavity $L/\xi\to 0$. Other parameters are $\mathcal{I}=0.95\,\mathcal{I}_{\text{cr}}$, $\xi=50\,\mu$m, $a_{s}=100\,a_{0}$, and $M=87\,u$. Inset: Three-dimensional BEC (in blue) placed at the cavity center aligned along the $x$-axis with a Gaussian mode profile of waist $\xi$. The broad pump laser propagates along the $y$-direction.
The formation of quantum droplets can be understood from analyzing $E_{0}$. For fixed $N$, (C1) and (C2) impose that the system exhibits an energy minimum with respect to the volume $V$ [41]. For repulsive contact interaction, the BEC mean-field energy Eq. (4) is positive and convex with respect to $V$. Conversely, due to $\Delta_{\rm C}<0$, which implies $\mathcal{I}<0$, the contribution of the cavity quantum fluctuations Eq. (7) turns out to be negative, which is typical for a roton mode. Thus, adding Eqs. (4) and (7) can lead to an energetic minimum as depicted in Fig. 1. To maintain the homogeneous phase, we need to stay below the Dicke phase transition. Since the latter occurs when the roton mode becomes soft, we derive the critical value $\mathcal{I}_{\text{cr}}=-2[(k^{2}/M)+2gn]/N(\sqrt{\pi}\operatorname{erf}[L/2\xi]\;\!\xi/L)^{4}$ from the condition $\Omega=0$. The emergence of $E_{\text{ac}}$ from a single roton mode has an important physical implication. In the thermodynamic limit, $N,L,\xi\to\infty$, while $N/V=$ const. and $L/\xi=$ const. Furthermore, the coupling of a single atom to the cavity vanishes, i.e., ${\cal G}_{0}\to 0$, so that $\mathcal{I}V=$ const. holds. This renders the roton’s energy contribution intensive, i.e., a finite-size effect. Consequently, in the limit of large $N$, the largest possible magnitude of $E_{\text{ac}}$, at $\Omega=0$, will eventually be dwarfed by any extensive energy term. Hence, self-bound droplets are only possible for systems of finite size.
To gain further understanding, we expand $V_{\rm C}(\bm{p},\bm{p}^{\prime})$ in Eq. (6) around $L/\xi=0$ to the second order. Restricting to $L/\xi<1$, this expansion is accurate for an experimental setup. The cavity quantum fluctuation correction is then expanded as $E_{\text{ac}}\approx E_{\text{ac}}^{(\infty)}+DL^{2}/2$. The first term $E_{\text{ac}}^{(\infty)}$ is the infinite interaction range energy correction, for $L\ll\xi$, indicated by the dashed horizontal line in Fig. 1. It does not depend on $L$ and is equivalent to taking the zeroth order approximation in $1/\xi$ of the Gaussian envelope of ${\cal G}(\bm{r})$. By this, we recover the known results for the infinite-range cavity-induced interaction. The second term, resulting from the specific shape of the envelope, contains the factor $D=-\mathcal{I}N/\big{[}\>\!12\xi^{2}\sqrt{1+(4gn+\mathcal{I}N)M/(2k^{2})}\;\!\big{]}$. We discard all contact interaction $gn\ll k^{2}/M$ in $E_{\text{ac}}$ which is only invalid around $\mathcal{I}\lesssim\mathcal{I}_{\text{cr}}$. Then this term has the form of a harmonic-like self-trapping potential as it is proportional to $L^{2}$. Therefore, the effective energy reads
$E_{0}(N,V)=E_{\text{ac}}^{(\infty)}+\frac{gN^{2}}{2V}-\frac{\mathcal{I}NV^{2/3}}{24\xi^{2}\sqrt{1+\mathcal{I}NM/(2k^{2})}},$ (8)
Figure 2: System pressure $P_{0}$ dependence on the atom number $N$ and the system size $L$ for $\mathcal{I}=0.95\,\mathcal{I}_{\text{cr}}$. The magenta line marks zero pressure. Inset: Droplet density (blue) and chemical potential (red) realized along this line. The absolute value of $\mathcal{I}$ decreases as $N$ increases. Remaining parameters as in Fig. 1.
As stated in (C1) and (C2), energy minimization with respect to the system size translates into a zero pressure condition $0=-(\partial E_{0}/\partial V)_{N}=P_{0}$ with positive bulk compressibility $0<K(P_{0}=0)=-V(\partial P_{0}/\partial V)_{N}|_{V=V_{0}}$, corresponding to a thermodynamically stable state [3]. The mean-field contact interaction yields a positive pressure $P_{\text{mf}}=gn^{2}/2$, while the roton mode corresponds to a negative pressure $P_{\text{ac}}=-D/(3L)$. As the cavity-induced interaction becomes stronger with increasing $\mathcal{I}$, the negative pressure $P_{\text{ac}}$ becomes comparable to the positive mean-field pressure $P_{\text{mf}}$. Their interplay results in a stable droplet with zero total pressure and positive bulk compressibility, as displayed in Fig. 1. The compressibility also obeys $K(P_{0}=0)/n<gn$. The corresponding sound velocity $c_{s}=\sqrt{K(P_{0}=0)/(Mn)}$ is modified by the roton mode contribution and turns out to be reduced with respect to the homogeneous counterpart $\sqrt{gn/M}$.
The resulting system pressure is depicted in Fig. 2 for varying particle number $N$ and system size $L$. The magenta line marks $P_{0}=0$, where the droplet size $L_{0}=V_{0}^{1/3}$ is determined by $N$ and other system parameters. We find negative pressure above this line and positive pressure below it, indicating the positive bulk compressibility $K(P_{0}=0)$ for the droplet corresponding to each atom number $N$. The inset shows the equilibrium density $n_{0}$ realized along the zero pressure line. As the number of particles increases, the droplet density decreases monotonically. The decrease is mainly caused by keeping the cavity interaction $\mathcal{I}=0.95\,\mathcal{I}_{\text{cr}}$ fixed due to $\mathcal{I}_{\text{cr}}\propto N^{-1}$. In principle, denser droplets can be realized by getting closer to the Dicke phase transition [39]. This is due to the divergence of the cavity pressure $P_{\text{ac}}$ when the roton mode vanishes as $\mathcal{I}\to\mathcal{I}_{\text{cr}}$. We note that the roton divergence goes along with the diverging condensate depletion into the roton mode. In line with contemporary experiments, we have chosen to stay $5\%$ below the critical point. With this we extract the leading order relation $V_{0}\sim N^{3/5}$ from our analytical expression for $V_{0}$.
Finally, we check whether the chemical potential remains negative [3] such that the droplet does not evaporate spontaneously, i.e., we analyze $[E_{0}(N-1)-E_{0}(N)]/N|_{P_{0}=0}>0$. The total chemical potential $\mu_{0}=gn+(\partial E_{\text{ac}}/\partial N)_{V}$ is composed of the positive mean-field and the negative roton contribution, respectively. At zero pressure, the cavity-induced term prevails over $gn$, resulting in a negative overall $\mu_{0}(P_{0}=0)<0$, thus avoiding self-evaporation [39]. As displayed in Fig. 2 (inset), the absolute value of the total chemical potential diminishes for growing $N$ in the same way as the density. Yet, it always remains negative.
Quantitative examination of the results in Figs. 1 and 2 reveals that, compared to typical BEC atom numbers and other droplet realizations [10, 11, 12, 13, 14], the droplet densities presented here are orders of magnitude more dilute. This roots in the BEC mean-field contact interaction that the cavity fluctuations have to overcome. In other droplet realizations, the mean field is almost completely suppressed from the outset. In addition, the homogeneous nature of the droplets is constrained by the Dicke phase transition. Droplets of any desired density can be obtained by approaching the Dicke critical point. However, due to technical limitations in experimental control, it is more viable to optimize other system parameters, such as, e.g., the equilibrium density which scales approximately as $n_{0}\sim g^{-3/5}$ with respect to the contact interaction strength [39]. It is also possible to optimize the setup with respect to the cavity mode width and explore the dependence $n_{0}\sim\xi^{-6/5}$ [39].
Although we have focused on the simplest form of a cavity-induced long-range interaction, the theory can account for any envelope function $f(\bm{r},\bm{r}^{\prime})$ and system dimension. However, analytical solvability requires the existence of an approximation of the envelope Fourier transform as in Eq. (3). Our theory, e.g., can easily be applied [39] to multimode cavities with an envelope $f(x,x^{\prime})=\exp[-(x-x^{\prime})^{2}/\xi^{2}]$ [34]. Those have been studied in a lattice with quantum Monte Carlo techniques [42] and in terms of numerical solutions for ring cavities [43].
@@ Results.
Conclusions. We have shown that a mean-field stable uniform BEC can be turned into a stable quantum droplet when coupled to a single-mode cavity with transverse pumping. The zero-point energy of the cavity quantum fluctuations provides a stabilizing attractive quantum correction to the mean-field BEC contact interaction energy for negative cavity detuning. The dependence of the attractive term on the BEC volume is crucially influenced by the shape of the cavity mode function, which determines the cavity-induced atom-atom interaction potential Eq. (2). The Fourier transform of the mode function envelope determines the roton eigenfrequency and thus the finite-size behavior of the roton energy contribution. An alternative view is that the cavity quantum fluctuations, once strongly coupled to the BEC, exert a significant negative pressure on the atomic cloud.
The equilibrium density of cavity-induced quantum droplets can be tuned by varying external cavity parameters, such as the strength of the transverse pump or the cavity detuning. The self-boundness should be the key experimental fingerprint of these quantum objects. If a BEC is prepared, coupled to the cavity, and then released by removing the confinement, the self-bound nature could be revealed by absorption imaging whether the cloud maintains its volume or expands. Finally, we note that the thermal pressure in the low-temperature limit $P_{\text{th}}=M^{3/2}\pi^{2}(k_{\rm B}T)^{4}/[\>\!90(gn)^{3/2}\>\!]$ would contribute in an experimental implementation. However, any realization would be based on the principal mechanism that the cavity quantum fluctuations realize competing interactions in the atoms that is fundamentally different from other quantum droplets and falls into the category (D3) in our classification.
@@
Acknowledgments.  We are grateful to Andreas Hemmerich and Hans Keßler for helpful discussions on the experimental parameters. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Collaborative Research Center SFB/TR185 (Project No. 277625399) (A.P.) and via the Research Grant No. 274978739 (M.R. and M.T.). We also acknowledge the support from the DFG Cluster of Excellence CUI: “Advanced Imaging of Matter” – EXC 2056 (Project ID 390715994).

### v2

@@ III.4 Thermal effects
IV Conclusions
@@
We have shown that a mean-field stable BEC can be turned into a stable quantum droplet when coupled to a single-mode cavity with transverse pumping. The zero-point energy of the cavity quantum fluctuations then provides an attractive quantum correction to the mean-field BEC contact interaction energy if the cavity is red detuned. The dependence of the attractive term on the BEC volume is crucially influenced by the shape of the cavity mode function, which determines the cavity-induced atom-atom interaction. The Fourier transform of the mode function envelope determines the roton eigenfrequency and thus the finite-size behavior of the roton energy contribution. An alternative view is that the cavity quantum fluctuations, once strongly coupled to the BEC, exert a significant negative pressure on the atomic cloud. It turns out that the spatial profiles of the light fields inducing the finite-range interaction are essential for droplet formation. The equilibrium density of cavity-induced quantum droplets can be controlled by varying the strength of the transverse pump or the cavity detuning. Self-boundedness is a key experimental hallmark of these quantum objects. If a condensate is prepared, coupled to the cavity, and then released by removing the external trapping along at least one axis, absorption imaging could reveal whether the cloud maintains its form or expands.
Although we have focused on the simplest form of a cavity-induced long-range interaction, the theory can account for any envelope function $f(\bm{r},\bm{r}^{\prime})$ and system dimension. Our theory, e.g., can directly be applied [36] to multimode cavities with an envelope $f(x,x^{\prime})=\exp[-(x-x^{\prime})^{2}/\xi^{2}]$ [31]. Those have been studied in a lattice with quantum Monte Carlo techniques [35] and in terms of numerical solutions for ring cavities [45].
Acknowledgments
@@
We are grateful to Andreas Hemmerich and Hans Keßler for helpful discussions on the experimental parameters. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Collaborative Research Center SFB/TR185 (Project No. 277625399) (A.P.) and via the Research Grant No. 274978739 (M.R. and M.T.). We also acknowledge the support from the DFG Cluster of Excellence CUI: “Advanced Imaging of Matter” – EXC 2056 (Project No. 390715994).
Appendix A Approximations leading to Eq. (3)
When the homogeneous mean-field approximation $\hat{\psi}(\bm{r})\to\langle\hat{\psi}(\bm{r})\rangle=\sqrt{n}$ is used in the second line of Eq. (1), one finds
$\frac{In^{2}}{32}\bigg{[}\int_{V}d^{3}\bm{r}\;\!\big{(}e^{ikx}+e^{-ikx}\big{)}\big{(}e^{iky}+e^{-iky}\big{)}\;\!e^{-(y^{2}+z^{2})/\xi^{2}}\bigg{]}^{2}\,.$ (9)
Since the atoms occupy a cube of volume $V=L^{3}$, it decomposes into the product of integrals given in Eq. (3). For discrete momenta $k=2\pi m/L$ for $m\in\mathbb{Z}$ and for the constraint $L<\xi$, the Gaussian $e^{-u^{2}/\xi^{2}}$ varies slowly compared to the fast oscillating exponential $e^{iku}$ for $k\neq 0$. Hence, we can approximate the Gaussian envelope by its spatial average
$\int_{-L/2}^{L/2}du\;\!e^{iku}e^{-u^{2}/\xi^{2}}$ $\approx\int_{-L/2}^{L/2}du\;\!e^{iku}\int_{-L/2}^{L/2}\frac{du}{L}e^{-u^{2}/\xi^{2}}$
$=\delta_{k0}\sqrt{\pi}\xi\operatorname{erf}\bigg{(}\frac{L}{2\xi}\bigg{)}\,.$ (10)
In this approximation, the spatial integral in Eq. (9) vanishes since $k\neq 0$.
Appendix B Bogoliubov theory
To include quantum fluctuations, we expand the Hamiltonian in Eq. (1) up to the second order in $\hat{\phi}(\bm{r})=\hat{\psi}(\bm{r})-\sqrt{n}$. The linear order vanishes upon applying Eq. (10). The quadratic order reads
$\hat{H}_{\rm qf}=$ $\int_{V}d^{3}\bm{r}\;\!\left[\hat{\phi}^{{\dagger}}\bigg{(}-\frac{\mbox{$\bm{\nabla}^{2}$}}{2M}+gn\bigg{)}\hat{\phi}+\frac{gn}{2}\Big{(}\hat{\phi}\hat{\phi}+\hat{\phi}^{{\dagger}}\hat{\phi}^{{\dagger}}\Big{)}\right]$
$+\frac{\mathcal{I}n}{2}\bigg{[}\int_{V}d^{3}\bm{r}\Big{(}\hat{\phi}+\hat{\phi}^{{\dagger}}\Big{)}\cos(kx)\cos(ky)\;\!e^{-\frac{y^{2}+z^{2}}{\xi^{2}}}\bigg{]}^{2}\,$, (11)
where we suppressed the spatial dependence of the fluctuation operators $\hat{\phi}$ and $\hat{\phi}^{{\dagger}}$ for brevity. The plane wave expansion of the fluctuation operator $\hat{\phi}(\bm{r})=\sum_{\bm{p}\neq\bm{0}}\hat{\phi}_{\bm{p}}e^{i\bm{p}\bm{r}}/\sqrt{V}$ leads to
$\hat{H}_{\rm qf}={}$ ${}\sum_{\bm{p},\bm{p}^{\prime}}{}^{{}^{\prime}}\hat{\phi}_{\bm{p}}^{{\dagger}}\bigg{(}\frac{\bm{p}^{\prime 2}}{2M}+gn\bigg{)}\hat{\phi}_{\bm{p}^{\prime}}\int_{V}\frac{d^{3}\bm{r}}{V}e^{-i(\bm{p}-\bm{p}^{\prime})\bm{r}}+\bigg{(}\frac{gn}{2}\hat{\phi}_{\bm{p}}\hat{\phi}_{\bm{p}^{\prime}}\int_{V}\frac{d^{3}\bm{r}}{V}e^{i(\bm{p}+\bm{p}^{\prime})\bm{r}}+h.c.\bigg{)}$
$+\frac{\mathcal{I}N}{2}\bigg{\{}\sum_{\bm{p}}{}^{{}^{\prime}}\bigg{[}\hat{\phi}_{\bm{p}}\int_{V}\frac{d^{3}\bm{r}}{V}\cos(kx)\cos(ky)e^{-i\bm{p}\bm{r}-\frac{y^{2}+z^{2}}{\xi^{2}}}+h.c.\bigg{]}\bigg{\}}^{2}\,$, (12)
where the primed sum indicates the omission of the zero momentum terms. The integrals are performed using the Kronecker integral representation $\int_{-L/2}^{L/2}du\;\!e^{ipu}=L\delta_{p0}$ and Eq. (10), yielding
$\hskip-5.69054pt\hat{H}_{\rm qf}$ $=\hskip-2.84526pt\sum_{\bm{p}}{}^{{}^{\prime}}\left[\bigg{(}\frac{\bm{p}^{2}}{2M}+gn\bigg{)}\hat{\phi}_{\bm{p}}^{{\dagger}}\hat{\phi}_{\bm{p}}+\frac{gn}{2}\Big{(}\hat{\phi}_{-\bm{p}}\hat{\phi}_{\bm{p}}+\hat{\phi}_{\bm{p}}^{{\dagger}}\hat{\phi}_{-\bm{p}}^{{\dagger}}\Big{)}\right]$
$\hskip-14.22636pt+\frac{\mathcal{I}N}{32V^{2}}\bigg{\{}\bigg{[}\sqrt{\pi}\xi\operatorname{erf}\bigg{(}\frac{L}{2\xi}\bigg{)}\bigg{]}^{2}\sum_{\bm{k}\in\mathcal{K}_{C}}\Big{(}\hat{\phi}_{\bm{k}}+\hat{\phi}_{\bm{k}}^{{\dagger}}\Big{)}\bigg{\}}^{2}\,$, (13)
with $\mathcal{K}_{\rm C}=\{({\!}\pm{\;\!\!}k\;\>\pm{\;\!\!}k\;\>0)^{T}\}$. Its eigenmodes are obtained by a standard Bogoliubov transformation to the quasiparticles of the modes unaffected by the cavity-induced interaction $\bm{p}\notin\mathcal{K}_{C}$. An additional transformation is required for those $\bm{k}\in\mathcal{K}_{C}$ that are coupled to the cavity. The former follow the standard Bogoliubov dispersion
$\omega_{\bm{p}}=\sqrt{\frac{\bm{p}^{2}}{2M}\bigg{(}\frac{\bm{p}^{2}}{2M}+2gn\bigg{)}}\,.$ (14)
To determine the eigenmodes for $\bm{k}\in\mathcal{K}_{C}$ it is convenient to express the Hamiltonian in terms of the quasi-position $\hat{x}_{\bm{p}}=\sqrt{M}(\hat{\phi}_{\bm{p}}+\hat{\phi}_{-\bm{p}}^{{\dagger}})/\sqrt{\bm{p}^{2}}$ and quasi-momentum $\hat{y}_{\bm{p}}=-i\sqrt{\bm{p}^{2}}(\hat{\phi}_{-\bm{p}}-\hat{\phi}_{\bm{p}}^{{\dagger}})/\sqrt{4M}$ operators of Ref. [46], yielding
$\hat{H}_{\rm qf}={}$ ${}\frac{1}{2}\sum_{\bm{p}\notin\mathcal{K}_{C}}{}^{{}^{\prime}}\bigg{(}\hat{y}_{\bm{p}}^{{\dagger}}\hat{y}_{\bm{p}}+\omega_{\bm{p}}^{2}\hat{x}_{\bm{p}}^{{\dagger}}\hat{x}_{\bm{p}}-\frac{\bm{p}^{2}}{2M}-gn\bigg{)}$
$+{}\frac{1}{2}\bigg{[}\hat{\vec{y}\;\!}^{{\dagger}}\mathbb{I}_{4\times 4}\>\!\hat{\vec{y}}+\hat{\vec{x}\;\!}^{{\dagger}}\underline{h}\;\!\hat{\vec{x}}-4\bigg{(}\frac{\bm{k}^{2}}{2M}+gn\bigg{)}\bigg{]}\,.$ (15)
Here we have defined $\hat{\vec{x}}=(\hat{x}_{\bm{k}_{1}}\;\>\hat{x}_{\bm{k}_{2}}\;\>\hat{x}_{\bm{k}_{3}}\;\>\hat{x}_{\bm{k}_{4}})^{T}$ and $\hat{\vec{y}}$ analogously, where $\mathcal{K}_{C}=\{\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},\bm{k}_{4}\}$. The cavity-mediated interaction in Eq. (2) only couples quasi-position operators so that the eigenmodes can be read off from the eigenvalues of the $4\times 4$ matrix coupling the four modes in $\mathcal{K}_{C}$
$\underline{h}=\omega_{\bm{k}}^{2}\;\!\mathbb{I}_{4\times 4}+\frac{\mathcal{I}N\bm{k}^{2}}{16M}\bigg{[}\frac{\sqrt{\pi}\xi}{L}\operatorname{erf}\bigg{(}\frac{L}{2\xi}\bigg{)}\bigg{]}^{4}\mathbb{J}_{4\times 4}\,$, (16)
where $\mathbb{J}_{4\times 4}$ is the matrix of all elements $1$. The eigenvalues are $\{\omega_{\bm{k}},\omega_{\bm{k}},\omega_{\bm{k}},\Omega\}$ with the first three obeying the standard Bogoliubov dispersion $\omega_{\bm{k}}$ from Eq. (14) and the last being affected by the cavity according to Eq. (5). Using these eigenmodes, we find the quantum fluctuations zero-point energy
$E_{\rm qf}=\frac{1}{2}\sum_{\bm{p}}{}^{{}^{\prime}}\bigg{(}\omega_{\bm{p}}-\frac{\bm{p}^{2}}{2M}-gn\bigg{)}+\frac{1}{2}\big{(}\Omega-\omega_{\bm{k}}\big{)}\,.$ (17)
The first term is the LHY correction of a Bose gas [9] that is to be evaluated in the continuum limit with the proper renormalization and equals $8VM^{3/2}(gn)^{5/2}/(15\pi^{2})$. In a weakly interacting dilute Bose gas, as studied here, the LHY correction is negligible compared to the mean-field energy in Eq. (4). The second term encapsulates the deviation of the roton dispersion from the Bogoliubov one due to the presence of the cavity-mediated interaction. It is exactly the cavity-induced quantum fluctuation correction $E_{\rm ac}$ given in Eq. (7).
Appendix C Derivation of the thermal pressure
At finite temperature, the effective energy $E_{0}(N,V,T)=E_{\rm mf}+E_{\rm ac}+E_{\rm th}$ acquires an additional thermal contribution [47, 44]
$\hskip-5.69054ptE_{\rm th}=\frac{1}{\beta}\sum_{\bm{p}}{}^{{}^{\prime}}\ln\big{(}1-e^{-\beta\omega_{\bm{p}}}\big{)}+\frac{1}{\beta}\ln\bigg{(}\frac{1-e^{-\beta\Omega}}{1-e^{-\beta\omega_{\bm{k}}}}\bigg{)}.$ (18)
The sum over the momenta is evaluated as an integral in the continuum limit. The ultra-low temperature regime implies large $\beta$, such that the Bogoliubov dispersion can be approximated as phononic, i.e., $\ln(1-e^{-\beta\omega_{\bm{p}}})\approx\ln(1-e^{-\beta\sqrt{gn\bm{p}^{2}/M}})$ [47]. This gives
$E_{\rm th}=-\frac{\pi^{2}VM^{3/2}}{90\beta^{4}(gn)^{3/2}}+\frac{1}{\beta}\ln\bigg{(}\frac{1-e^{-\beta\Omega}}{1-e^{-\beta\omega_{\bm{k}}}}\bigg{)}\,.$ (19)
The resulting thermal pressure is then
$P_{\rm th}=\frac{\pi^{2}M^{3/2}}{36\beta^{4}(gn)^{3/2}}\,$, (20)
where the remaining roton contribution is negligible.